OpenStudy (anonymous):

what is the exact value of cos 15 in radical form? the caculator doesn't show it in radicals..

OpenStudy (sirm3d):

method 1: use the identity \[\cos(A-B)=\cos A \cos B + \sin A \sin B\] on \[\cos 15^0=\cos(45^0-30^0),\;A=45^0,B=30^0\] method 2: use the identity \[\cos(A/2)=\sqrt{\frac{1+\cos A}{2}}\] on \[\cos 15^0=\cos(30^0/2),\;A=30^0\]